### v2

# A Comprehensive Survey on Deep Multimodal Learning with Missing Modality

Abstract.
During multimodal model training and reasoning, data samples may miss certain modalities and lead to compromised model performance due to sensor limitations, cost constraints, privacy concerns, data loss, and temporal and spatial factors. This survey provides an overview of recent progress in Multimodal Learning with Missing Modality (MLMM), focusing on deep learning techniques. It is the first comprehensive survey that covers the historical background and the distinction between MLMM and standard multimodal learning setups, followed by a detailed analysis of current MLMM methods, applications, and datasets, concluding with a discussion about challenges and potential future directions in the field.
multimodal learning, deep learning, representation learning, multimodal fusion, audio signal processing, multimedia, missing modality, imperfect modality, incomplete modality
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learningccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing
1. Introduction
(a) Multimodal Model Description
(b) Full-Modality Samples vs. Missing-Modality Samples
Figure 1. (a) AI models can input and process data in a variety of different modalities; (b) Description of a three-modality scenario with full- and missing-modality samples. We abbreviate all “Modality” as “Mod” in all figures of this paper and visualize missing modalities with nearly transparent colors and dashed boxes.
Multimodal learning has become a crucial field in Artificial Intelligence (AI). It focuses on integrating and analyzing various data types, including visual, textual, auditory, and sensory information (Figure 1(a)). This approach mirrors the human capacity to combine multiple senses for better understanding and interaction with the environment. Modern multimodal models leverage the robust generalization capabilities of deep learning to uncover complex patterns and relationships that uni-modality systems might not detect. This capability is advancing work across multiple domains, including computer vision. Recent surveys in these fields highlight the significant impact of multimodal approaches, demonstrating their ability to enhance performance and enable more sophisticated AI applications (Baltrušaitis et al., 2018; Zhao et al., 2024).
However, multimodal systems often face the challenge of missing or incomplete data in real-world applications. This occurs due to various factors such as sensor malfunctions, hardware limitations, privacy concerns, environmental interference, and data transmission issues. As illustrated in  Figure 1(b), in a three-modality scenario, data samples can be categorized as either full-modality (containing information from all three modalities) or missing-modality (lacking data from one or more modalities entirely). These problems can arise at any stage from data collection to deployment, significantly impacting model performance. Real-world examples of this problem are prevalent across various domains. In affective computing, researchers (Cohen et al., 2004; Sebe et al., 2005) have encountered samples with only usable images or audio due to camera obstructions or excessive microphone noise. Similarly, in space exploration, NASA’s Ingenuity Mars helicopter (Donaldson, 2024) faced a missing modality challenge when its inclinometer failed due to extreme temperature cycles on Mars. To address this issue, NASA applied a software patch that modified the initialization of navigation algorithms (Team, 2022). In the field of medical AI, privacy concerns often result in certain modalities being unavailable for some data samples, creating inherent missing modalities in multimodal datasets (Zhang et al., 2023c). The unpredictability of real-world scenarios and diversity of data sources further compound this challenge. Consequently, developing robust multimodal systems that can perform effectively with missing modalities has become a crucial focus in the field.
In this survey, we refer to the challenge of handling missing modalities as the ”missing modality problem.” We name solutions to this problem Multimodal Learning with Missing Modality (MLMM). This approach contrasts with the traditional setup that utilizes full modalities, which we name Multimodal Learning with Full Modality (MLFM). Specifically, in MLFM tasks, given a $H$-modality dataset, we are typically required to train a model can process and fuse information from all $H$ modalities to make predictions. During training and testing, samples with complete information from all $H$ modalities are used. In contrast, MLMM tasks use fewer than $H$ modalities during training or testing due to factors such as data collection limitations or constraints in deployment environments. The primary challenge in MLMM lies in dynamically and robustly handling and fusing information from any number of available modalities during training and testing, while maintaining performance comparable to that achieved with full-modality samples.
Figure 2. Our Methodology Taxonomy of Deep Multimodal Learning with Missing Modality Methods from four dimensions: modality augmentation, feature space engineering, architecture engineering, and model selection. “MLLMs” means multimodal large language models.
This survey encompasses recent advances in MLMM and its applications across various domains, including information retrieval, remote sensing, and robotic vision. We provide a fine-grained taxonomy of MLMM methodologies, application scenarios, and corresponding datasets. Our work expands upon existing MLMM surveys that focus on specific domains such as medical diagnosis (Zhou et al., 2023; Shah et al., 2023; Azad et al., 2022), sentiment analysis (Wagner et al., 2011), and multi-view clustering (Chao et al., 2021). By offering a comprehensive overview of current research and identifying promising directions for future work, this survey aims to contribute to the development of more robust and adaptable multimodal learning systems. These advancements are crucial for deploying intelligent systems in challenging environments, ranging from the harsh conditions of planetary exploration to the dynamic and unpredictable settings of everyday life. Main contributions of this survey are three-fold:
  1. (1)
    A comprehensive survey of MLMM applications across diverse domains, accompanied by an extensive compilation of relevant datasets, highlighting the versatility of MLMM in addressing real-world challenges.
  2. (2)
    A novel, fine-grained taxonomy of MLMM methodologies, with a multi-faceted categorization framework based on recovery strategies, integration stages, parameter-efficiency approaches, and attention mechanisms.
  3. (3)
    An in-depth analysis of current MLMM approaches, their challenges, and future research directions, contextualized within the proposed taxonomical framework.
Paper Collection: In our literature search methodology, we primarily source papers from Google Scholar and major conferences/journals in Artificial Intelligence, Machine Learning, Computer Vision, Natural Language Processing, Audio Signal Processing, Data Mining, Multimedia, Medical Imaging, and Remote Sensing etc. We have compiled a total of 315 significant papers from the period spanning 2012 to 2024 September. Our search strategy involves using keywords such as “incomplete”, “missing”, “partial”, “absent”, and “imperfect”, coupled with terms like “multimodal learning”, “deep learning”, “representation learning”, and “neural network”.
Survey Organization: Firstly, we explain the rough history of deep MLMM, background and motivation of this survey in Section 1. In Section 2, we categorize the methods of existing deep MLMM solutions from a methodological perspective, detailing them in four dimensions (Figure 2). We then summarize the current application scenarios and corresponding datasets used in this field in Section 3. In Section 4, we discuss the unresolved challenges and future directions of this field. Finally, in Section 5, we present our conclusions drawn from the exploration of deep MLMM.
2. Methodology
We classify and discuss existing deep MLMM methods based on our methodology taxonomy from four major dimensions: modality augmentation, feature space engineering, architecture engineering, and model selection.
2.1. Modality Augmentation
We categorize modality augmentation methods addressing the missing modality problem at the original data of modality level into two types. The first is the modality composition method, which uses zero/random values, directly copied data from similar instances, or matching samples obtained via retrieval algorithms, composited with the missing-modality sample to form a full-modality sample. The second is the modality generation method, which generates the original data of the missing modality using generative models such as Auto Encoders (AEs) (Hinton and Zemel, 1993), Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), or Diffusion Models (Ho et al., 2020).
Figure 3. Zero/Random values composition method. If we assume modality 2 is missing, then this modality will be replace with zero/random values. “DNN” in all figures of this paper means different kinds of Deep Neural Networks.
Figure 4. General idea of retrieval-based modality composition methods. We retrieval one or more samples by randomly selecting or using simple retrieval algorithms like KNN and its variants from past or same-category samples that have required missing modalities, and then compose them with the input missing-modality sample to form a full-modality sample.
2.1.1. Modality Composition Method
Removing missing-modality samples is common and simple as a multimodal dataset preprocessing method when the dataset has few missing-modality samples. But this method will reduce the dataset size and make some rare samples disappear when the dataset contains many missing-modality samples. The composition method is widely employed for its simplicity and effectiveness and also keep the dataset size. One typical method is the zero/random values composition method which replaces the missing modality data with zero/random values, as shown in Figure 3. In recent research (Chen et al., 2020; Sun et al., 2024a; Liu et al., 2023c; Malitesta et al., 2024), they are often used as baselines for comparison with proposed methods. About the missing sequential data problem, such as missing frames in videos, the similar Frame-Zero method (Parthasarathy and Sundaram, 2020) was proposed to replace the missing frames and composed with available frames. These methods are prevalent in typical multimodal learning training procedures. By using this, multimodal models can balance and integrate information from different modalities when making predictions. This helps prevent the model from relying too heavily on one or a few modalities to some extend, thereby enhancing the robustness. Additionally, these methods fail to generalize well when most samples in the dataset are the missing-modality samples. While, other composition method is based on retrieval algorithms (Figure 3), which involves filling in the original data of the missing modality by copying/averaging the original data from retrieved samples that have same modality and category. Some convenient methods randomly select a sample that has same category and required missing modality from other samples. The selected sample is then composed with the input missing-modality to form a full-modality sample for training. For example, researchers (Yang et al., 2024) proposed Modal-mixup to complete the training datasets by randomly complementing same-category samples with missing modalities. However, such methods cannot solve the missing modality problem during the testing phase. Regarding missing frames in streaming data such as videos, researchers proposed using Frame-Repeat (Parthasarathy and Sundaram, 2020) to make up for the missing frames by using past frames. Some work (Yang et al., 2024; Campos et al., 2015; Gong et al., 2021) also has attempted to use K-Nearest Neighbors (KNN) or its variants that can measure the similarity between samples to retrieve the best-matched samples for composition. For these matched samples, we can select the sample with the highest score or obtain the average values of these samples to supplement the missing modality data. Their experiments have shown that KNN-based methods generally perform better than above methods such as random selection and can handle missing modality during testing. But straightforward simple clustering often suffers from high computational complexity, sensitivity to imbalanced data, and high memory overhead. Also, retrieval-based modality combination methods are not applicable to pixel-level tasks such as segmentation and are only suitable for simple tasks (e.g., classification) because they may lead to model confusion. In addition, above all methods can complete datasets with missing modalities, it reduces the diversity of the dataset. This is especially problematic with datasets having a high modality missing rate (means most samples are missing-modality samples), as it increases the risk of overfitting to certain categories with few full-modality samples.
(a) Individual Modality Generation Method
(b) Unified Modality Generation Method
Figure 5. Description of two typical Modality generation methods. We set modality 2 as the missing modality for examples and use other available modalities to generate modality 2. “GEN” in both figures represents generative networks. (a) We set up a modality-2 generator (GEN-2) from other modalities. In the actual training process, modality 1 and 3 each have such a GEN. For simplicity, we omit the two GENs. (b) All modalities are input and generated together by using a single GEN.
2.1.2. Modality Generation Method
Various matrix imputation methods (Gong et al., 2021) were proposed in missing data research, leveraging latent correlations among matrix elements. However, missing data often appears in large blocks in multimodal datasets, making conventional methods inefficient for large-scale processing and high-dimensional computations. With deep learning, synthesizing missing modalities has become more efficient. Current methods for generating original data of missing modalities are divided into individual and unified generative methods. Individual modality generation method trains an individual generative model for each modality in case of any missing modality situation, which is shown in Figure 5(a). Early works used Gaussian processes (Mario Christoudias et al., 2010) or Boltzmann machines (Sohn et al., 2014) to generate missing modalities from available inputs. With deep learning, methods like AEs and U-Net(Ronneberger et al., 2015) can be used for original modality data generation. Li et al.(Li et al., 2014) used 3D-CNN to generate positron emission tomography (PET) data from magnetic resonance imaging (MRI) inputs. Chen et al.(Chen et al., 2024b) addressed missing modalities in MRI segmentation by training U-Net models to generate other two modalities from MRI. Recent work (Ma et al., 2021a) used AEs as one of the baselines to complete datasets by training one AE per modality. In domain adaptation, Zhang et al. (Zhang et al., 2021) proposed a Multi-Modality Data Generation module with domain adversarial learning to generate each missing modality by learning domain-invariant features. GANs significantly improved image generation quality by using a generator to create realistic data and a discriminator to distinguish it from real data. Researchers began replacing AEs and U-Nets with GANs for missing modality generation. For example, GANs generated original data of missing modalities using latent representations of existing ones in breast cancer prediction (Arya and Saha, 2021), and WGANs were applied in sentiment analysis (Wang et al., 2022). In remote sensing, Bischke et al.(Bischke et al., 2018) used GANs to generate depth data, improving segmentation over RGB-only models. GANs were also used in robotic recognition to generate RGB and depth images(Gunasekar et al., 2020). Recent studies (Ma et al., 2021a) show that GANs outperform AEs in generating more realistic missing modalities and can lead to better downstream-task model performance. Recently, the introduction of Diffusion models has further improved image generation quality. Wang et al. proposed the IMDer method (Wang et al., 2024a), which uses available modalities as conditions to help diffusion models generate missing modalities. Experiments showed diffusion reduces semantic ambiguity between recovered and missing modalities and achieves good generalization performance than previous works. But, training an individual generator for each modality is inefficient and time-consuming, failing to capture the latent correlations between modalities. Researchers develop another type of generative method, the unified modality generation method, which trains a unified model that can generate all modalities simultaneously (Figure 5(b)). One representative model is Cascade AE (Tran et al., 2017), which stacks AEs to capture the differences between missing and existing modalities for generating all missing modalities. Recent researchers, such as Zhang et al. (Zhang et al., 2023b), have attempted to use attention mechanisms and max-pooling layers to integrate features of existing modalities, enabling modality-specific decoders to generate each missing modality. Experiments demonstrated that this method is more effective than using max-pooling alone (Chartsias et al., 2017) to integrate features from any number of modalities for generating missing modalities. The above methods for generating original data for missing modalities can mitigate performance degradation to some extent. However, training a generator that can produce high-quality, real-world-distribution-like missing modalities remains challenging, especially when the training dataset contains a few full-modality samples. Additionally, the modality generation model significantly increases storage requirements. As the number of modalities grows, the complexity of these generative models also increases, further complicating the training process and resource demands.
Figure 6. Illustration of representation constraint methods (include regularization-based and correlation driven methods). Set modality 2 as the missing modality. Between the features of different modalities, correlation or regularization terms are used.
2.2. Feature Space Engineering
The following introduces methods to address the missing modality problem at the feature space. First, we present two constraint-based approaches that enhance the learning of more discriminative and robust representations by imposing specific constraints (Figure 6). One method involves regularization to improve the effectiveness and generalization of the learned representations. Another method focuses on maximizing correlations, using specific measurements to strengthen the relationships between features. Next, the representation composition method, which can borrow the solutions discussed in Section 2.1.1, operates at the feature level of the modalities or employs arithmetic operations to handle a dynamic number of modalities. Finally, we introduce the representation generation method, which can generate the representations of missing modalities.
2.2.1. Regularization Based Method
There are few regularization based methods, such as Tensor Rank Regularization (Liang et al., 2019), which introduces tensor rank minimization as a regularization term in the training process. This method can automatically handle missing or noisy data without prior knowledge of data imperfections. It combines the temporal nonlinear transformation of multimodal data with simple regularization techniques of tensor structures, helping to learn true correlations and latent structures. This effectively mitigates the imperfection of input data and addresses the missing modality problem. Due to the limited number of this kind of methods, they have only been verified on time series data, and whether they can work in other fields requires extensive verification.
2.2.2. Correlation Driven Method
Some methods train models by coordinating the correlation between different modal features. For example, Wang et al. (Wang et al., 2020a) use a Deep Canonical Correlation Analysis (CCA) module to maximize feature associations within the same category, enabling training on incomplete datasets. Ma et al. (Ma et al., 2021b) proposed a Maximum likelihood function to characterize conditional distributions of full-modality and missing-modality samples during training. Other researchers (Liu et al., 2021a) have added constraints based on the Hilbert-Schmidt Independence Criterion to help the model learn how to complete missing modality features from full modality features. These methods all aid models in learning how to complete missing modality features or training on incomplete datasets by learning the similarity or correlation between features. A drawback of above methods is that they perform well only when two or three modalities are used as input. Even when trained on a dataset with missing modalities, some methods still struggle to effectively address the missing modality problem during testing. Additionally, as the number of modalities increases, the variety of trade-offs in loss functions becomes more complex, making it challenging to achieve good generalization performance.
Figure 7. The general idea of the arithmetic method. The representation of any number of modalities can be combined to a fixed dimension that is the same as the dimension required by the subsequent DNN layer.
2.2.3. Representation Composition Method
There are two types of representation composition methods: one is the retrieval-based composition method, which attempts to recover the missing modality representation by borrowing existing samples or modality information, similar to the original data composition method in Section 2.1.1. They typically use pre-trained feature extractors to generate features from available samples, storing them in a feature pool (Zhi et al., 2024; Sun et al., 2024a). Cosine similarity is then used to retrieve matched features for the input sample, with the average of these missing-modality features in top-k samples filling the representation of missing modalities. Additionally, some methods, such as Missing Modality Feature Generation (Wang et al., 2023c), replace missing modality features by averaging the representations of available modalities, assuming similar feature distributions across modalities. Another representation composition method, the arithmetic method, dynamically processes modality representations without using learnable parameters like pooling layers (Figure 7). Researchers (Zhou et al., 2021; Delgado-Escano et al., 2021; Marín-Jiménez et al., 2021; Wang et al., 2023c) have fused features through operations like average/max pooling or addition, offering low computational complexity and efficiency. However, drawbacks include potential loss or dilution of important information. To address this, merge operations (Delgado-Escano et al., 2021; Marín-Jiménez et al., 2021) use a sign-max function, which selects the highest absolute value from feature vectors, yielding better results by preserving both positive and negative activation values. TMFormer (Zhang et al., 2024) introduces token merging based on cosine similarity, while others calculate moments of available modalities for feature fusion through addition and concatenation (Havaei et al., 2016). Similar approaches (Zheng et al., 2024a, b; Zhang et al., 2024) focus on selecting key vectors or merging tokens to handle missing modalities. Recent works (Zheng et al., 2024a, b) focus on selecting key features or vectors for handling missing modalities. Despite these advancements, methods can still struggle if a major modality is missing. Therefore, those representation composition approaches not only allow the model to flexibly handle features from any number of modalities but also enable learning without/few learnable parameters. It is also difficult to capture the inter-modality relationships through the learning way, as methods like selecting the largest vector or the feature with the highest score is difficult to fully represent the characteristics of all modalities.
(a) Indirect-to-Task Representation Generation Method
(b) Direct-to-Task Representation Generation Method
Figure 8. Description of two typical representation generation methods. We assume modality 2 is missing here. (a) This kind of methods are supervised by the loss of two tasks, auxiliary reconstruction and downstream tasks. Note that these two tasks sometimes can be trained separately, first training the reconstruction task, discarding the generator, and then training the downstream task, such as the training progress of ActionMAE (Woo et al., 2023). While (b) methods directly use a generative network (“GEN-2” in figure) to generate representations of the missing modality 2 during training and inference without reconstructing the original data of modality 2.
2.2.4. Representation Generation Method
Compared to original data generation methods in Section 2.1.2, hierarchical representation generation can integrate seamlessly with existing multimodal frameworks. Current methods fall into two categories: (1) Indirect-to-task representation generation methods (Figure 8(a)) treat modality reconstruction as an auxiliary task during training, helping the model intrinsically generate missing modality representations for downstream tasks. Since the auxiliary task aids in representation generation during training but is dropped during inference of downstream task, it is termed “indirect-to-task.” (2) Direct-to-task representation generation methods (Figure 8(b)) train a small generative model to directly map available modality information to the representation of missing modalities.
Indirect-to-Task Representation Generation Methods: A common architecture for indirectly generating missing modality representations is the “encoder-decoder” model. During training, modality-specific encoders extract features from available modalities, one/some reconstruction decoders reconstructs the missing modality, and one downstream module supervised by downstream task loss for prediction. Those reconstruction decoders are discarded during inference, where predictions rely on both the generated missing modality representations and the existing ones. This approach typically employs Multi-Task Learning, training both downstream tasks and auxiliary tasks (modality reconstruction) simultaneously. Some methods, inspired by Masked Autoencoder (He et al., 2022), split reconstruction and downstream task training. Based on the input features used by reconstruction decoders, we divide these methods into two categories. The first category, Before Fusion methods, inputs the features from modality-specific encoders directly into each specific reconstruction decoder for missing modality reconstruction. For example, MGP-VAE (Hamghalam et al., 2021) uses the VAE encoder’s output for Gaussian Process prediction and reconstruction decoders. Similarly, PFNet (Zheng et al., 2021) employs RGB and thermal-infrared modalities for pedestrian re-identification, feeding outputs of each encoder into corresponding decoders before fusing them together. Li et al. (Li et al., 2024b) adopted a similar approach for brain tumor segmentation, reconstructing each modality before fusing encoder outputs. The second category, After Fusion methods, uses the fused features from modality-specific encoders as the input of reconstruction decoder. Tsai et al.(Tsai et al., 2018) introduced a Multimodal Factorization Model where modality-specific encoder-decoders reconstruct the missing modality, while the separate downstream task model employs a multimodal encoder-decoder. To recover the the missing modality features, the outputs of the downstream task model’s encoder are not only fused with the outputs of each modality-specific encoder for inputting to the each corresponding reconstruction decoder, but also used for predictions of the downstream task decoder. Chen et al.(Chen et al., 2019) also decoupled the encoders for reconstruction and downstream tasks, showing that disentangling features helped eliminate modality-specific information. Jeong et al.(Jeong et al., 2022) proposed a simpler approach using multi-level skip connections in a U-Net architecture to fuse features for missing modality reconstruction, achieving better generalization in brain tumor segmentation than Chen et al. Further, Zhou et al.(Zhou, 2023) and Sun et al. (Sun et al., 2024b) demonstrated effective cross-modal fusion techniques to enhance reconstruction and segmentation. Most of these methods train reconstruction and downstream tasks concurrently. However, some approaches, such as ActionMAE (Woo et al., 2023) and M3AE (Liu et al., 2023c), first train reconstruction tasks and then fine-tune the pretrained encoders for downstream tasks. Despite achieving strong generalization, these methods are limited by the requirement of supervising by the original data of missing modalities.
Direct-to-Task Representation Generation Methods: In contrast to indirect methods, these approaches directly map information from existing modalities to the missing modality representation. The general concept is illustrated in Figure 8(b). Early work (Mario Christoudias et al., 2010) using Gaussian processes to “hallucinate” missing modalities inspired Hoffman et al. (Hoffman et al., 2016) to propose a Hallucination Network (HN) that predicts missing depth modality features from RGB images. The HN (student) aligns intermediate layer features with a teacher network (trained on depth images) using an L2 loss, allowing RGB inputs to generate depth features to address missing modality challenges. This method has been extended to pose estimation (Choi et al., 2017) and land cover classification (Kampffmeyer et al., 2018), where the HN generates heat distribution and depth features. In tasks like sentiment analysis, a translation module can map representations of available modalities to the missing ones, with L2 loss applied to align outputs with those from networks trained on missing modalities (Huan et al., 2023). A notable recent work, SMIL (Ma et al., 2021a), employs Bayesian learning and meta-regularization to directly generate missing modality representations. SMIL has shown strong generalization on datasets with a high rate of missing modality samples, handling up to 90% missing-modality samples. Subsequent work (Li et al., 2022c) addressed challenges such as distribution inconsistency, failure to capture specific modality information, and lack of correspondence between generated and existing modalities. They introduced solutions such as discriminator-based consistency loss, memory representativeness, and structural alignment. Researchers have also explored cross-modal distribution transformation methods to align distributions of generated and existing modality representations, enhancing discriminative ability (Wang et al., 2023b). In audio-visual question answering (AVQA), Park et al. (Park et al., 2024) proposed a Relation-aware Missing Modal Generator to generate pseudo features of missing modalities, improving robustness and accuracy. Compared to indirect methods, direct generation avoids the cumbersome reconstruction of raw data, using a feature generator that is easier to train and integrate into multimodal models. Recently, Kim et al.(Kim and Kim, 2024) combined self-supervised learning, Joint Embedding Predictive Architecture, and parameter-efficient learning to help the model generate missing modality features. Additionally, Guo et al.(Guo et al., 2024) proposed a Missing Modality Generation Module for prompt learning, mapping prompts of available modalities to those of missing ones. Lin et al. (Lin et al., 2024) introduced U-Adapter, which performs cross-modal fusion while preventing unreliable information due to domain shifts, effectively solving the missing modality problem.
In summary, above representation generation methods are all limited by the imbalance of missing modalities in the training dataset or the limited number of full-modality samples, which may lead to the possibility of being dominated by existing modalities. In addition, since indirect-to-task representation generation methods can access the missing modality data of the missing modality samples during training, which means that the training dataset is actually complete, this type of method usually performs better than direct-to-task representation generation methods.
2.3. Architecture Engineering
Different from above methods of generating modality or representations of modalities, some researchers adjust the model architecture to adapt to the missing-modality cases. We divide them into four categories according to their core contributions in dealing with missing modalities: attention-based methods, distillation-based methods, graph learning-based methods, and multimodal large language models (MLLMs).
(a) Intra-Modality Attention Method
(b) Inter-Modality Attention Method
Figure 9. General idea of two typical attention fusion methods. We set modality 2 as the missing modality. In (a), the missing modality 2 in intra-modality attention (“Intra-Atten” in figure) can be skipped. Since each Intra-Atten of different modalities shares the fusion representation (like learnable query in (Ge et al., 2023) or bottleneck tokens in (Nagrani et al., 2021)), after calculating them one by one, the shared learnable representation can be regarded as the fusion of representations of available modalities. In (b), “Missing Modality Mask” is the custom attention mask that is generated according to the missing modality 2 and can help inter-modality attention (“Inter-Atten” in figure) ignore the missing modality 2. By setting the tokens of modality 2 to zero or negative infinity through masking, we can force the attention mechanism to ignore the missing modality 2.
2.3.1. Attention Based Method
In the self-attention mechanism (Vaswani et al., 2017), each input is linearly transformed to generate Query, Key, and Value vectors. Attention weights are computed by multiplying the query of each element with the keys of others, followed by scaling and softmax to ensure the weights sum to 1. Finally, a weighted sum of the values generates the output. We classify attention-based MLMM methods into two categories: attention fusion methods, which integrate multimodal information (e.g., text, vision, audio) and are suitable for various networks; and transformer-based methods, which stack attention layers to handle large-scale data with global information capture and parallelization.
Attention Fusion Based Methods: Due to the powerful ability of attention to capture key features and its plug-and-play benefits, many methods have adopted attention mechanisms. We categorize them into two types: intra- and inter-modality attention methods. Intra-modality attention methods compute attention for each modality independently before fusing them, as shown in Figure 9(a). This approach focuses on relationships within a single modality (e.g., text, image, or audio), and the fusion between modalities is achieved by learnable fusion representations. For instance, in 3D detection for autonomous vehicles, the BEV-Evolving Decoder (Ge et al., 2023) handles sensor failures by sharing same BEV-query with each modality-specific attention modules, allowing fusion of any number of modalities. Similarly, in clinical diagnosis, Lee et al. (Lee et al., 2023a) proposed modality-aware attention to perform intra-modality attention and predict decisions by fusing bottleneck fusion tokens (Nagrani et al., 2021), addressing learning with missing modalities. Inter-modality attention methods, often based on masked attention, treat missing modality features as masked vectors to better capture dependencies across modalities, as illustrated in Figure 9(b). Unlike conventional cross-modal attention, masked attention models share the same parameters across all embeddings, allowing flexible handling of missing modalities. For example, Qian et al.(Qian and Wang, 2023) designed an attention mask matrix to ignore missing modalities, improving model robustness. Similarly, DrFuse(Yao et al., 2024) decouples modalities into specific and shared representations, using the shared ones to replace missing modalities, with a custom mask matrix to help model ignore the specific representations of the missing modality.
Figure 10. Description of general idea of joint representation learning methods. Modality 2 is set as the missing modality. The missing modality 2 tokens in can be replaced by learnable tokens or specific tokens. The modality encoders (red dotted rectangle) can be replaced by linear projection layers in some methods.
Transformer Based Methods: We divide transformer-based methods into two methods: joint representation learning (JRL) and parameter efficient learning (PEL) according to full parameter training and a small amount of parameter fine-tuning. By utilizing stacked multimodal transformer layers with attention, the multimodal transformer can learn joint representations from any number of modality tokens (Figure 10). Gong et al. (Gong et al., 2023) introduced an egocentric multimodal task, proposing a transformer-based fusion module with a flexible number of modality tokens and a cross-modal contrast alignment loss to map features into a common space. Similarly, Mordacq et al. (Mordacq et al., 2024) leveraged a Masked Multimodal Transformer, treating missing modalities as masked tokens for robust joint representation learning. Ma et al. (Ma et al., 2022) proposed an optimal fusion strategy search strategy to enhance the performance of multimodal transformers in handling missing modalities. Radosavovic et al.(Radosavovic et al., 2024) introduced mask tokens [M] for an autoregressive transformer to manage missing modalities, successfully deploying the model in real-world scenarios. With the rise of pre-trained transformer models, PEL methods have been developed to fine-tune these models by training few parameters, leveraging their generalization abilities for various tasks. Given their focus on pre-trained models and limited parameter updates, we categorize PEL methods separately, despite some overlap with above JRL methods. Two common fine-tuning methods for pre-trained models are prompt and adapter tuning. Initially used in natural language processing, prompt tuning optimizes input prompts while keeping model parameters fixed. It has since been extended to multimodal models, optimizing multimodal prompt embedding to handle diverse inputs like text and images. Jang et al.(Jang et al., 2024) introduced modality-specific prompts to address limitations in earlier methods, which merge when all modalities are present and allow the model to update all learnable prompts during training. Liu et al.(Liu et al., 2024b) further improved this by proposing Fourier Prompt, which uses fast Fourier transform to encode global spectral information into learnable prompt tokens which can be used to supplement missing-modality features, enabling cross-attention with feature tokens to address missing modalities. Adapter tuning, on the other hand, involves inserting lightweight adapter layers into pre-trained models to adapt to new tasks without modifying the original parameters. Qiu et al. (Qiu et al., 2023) proposed a method that uses a classifier to identify missing-modality cases and the intermediate features as the missing-modality prompt to cooperate with a lightweight Adapter to address missing modality problems for brain tumor segmentation.
Although attention-based fusion mechanisms in above methods can effectively help deal with the missing modality problem in any framework, none of them cares about the missing modality that may contain important information required for prediction. In the transformer-based methods, although the PEL method has a very fast fine-tuning speed and requires very few computing resources, its performance is still not comparable to that of JRL. But the JRL method is limited by a large amount of computing resources and requires relatively large datasets to achieve good performance.
Figure 11. Description of representation distillation methods. Here we set modality 2 is missing. In order to distinguish the representations of teachers and students, we add “-T” and “-S” after their representations in the figure. We refer to intermediate distillation as using features from any combination of intermediate layers within models.
(a) Mean Teacher Distillation Method
(b) Self Distillation Method
Figure 12. Description of two process-based distillation methods. (a) “EMA” means exponential moving average. To demonstrate the general architecture of Mean Teacher Distillation, we set the loss on the logits representation. (b) Self-distillation methods involve models acting as both the teachers and students, where most of them uses their own predictions from each branch to refine their representations during training.
2.3.2. Distillation Based Method
Knowledge distillation transfers knowledge from a teacher model to a student model (Hinton et al., 2015). The teacher model, with access to more information, helps the student reconstruct missing modalities. Below, we categorize two types of distillation methods for addressing this problem. First is the representation-based distillation, this method transfers rich representations from the teacher model to help the student capture and reconstruct missing modality features. We classify approaches based on whether they use decision-level or intermediate features. Figure 11 illustrates the procedures. Response distillation methods focus on transferring the teacher model’s logits to the student, helping it mimic probability distributions. Wang et al. (Wang et al., 2020b) trained modality-specific teachers for missing modalities, then used their soft labels to guide a multimodal student. Hafner and Ban (Hafner and Ban, 2023) employed logits from a teacher model trained with optical data to supervise a reconstruction network for approximating missing optical features from radar data. Pramit et al. (Saha et al., 2024) proposed a Modality-Aware Distillation, leveraging both global and local teacher knowledge in a federated learning setting to handle missing modalities. Orzikulova et al. (Orzikulova et al., 2024) used logits for privacy-preserving knowledge transfer in missing modality scenarios. Intermediate distillation methods align intermediate features between teacher and student models. Shen et al. (Shen and Gao, 2019) used Domain Adversarial Similarity Loss to align intermediate layers of teacher and student, improving segmentation in missing modality settings. Zhang et al.(Zhang et al., 2023c) applied intermediate distillation in endometriosis diagnosis by distilling features from a TVUS-trained teacher to a student using MRI data. Then is the process-based method, which focuses on overall distillation strategies, like Mean Teacher Distillation (MTD) (Tarvainen and Valpola, 2017) and self-distillation(Zhang et al., 2019). These methods emphasize procedural learning over direct representation transfer. MTD enhances stability by using the exponential moving average of the student model’s parameters as a teacher (Figure 12(a)). Chen et al. (Chen et al., 2024a) applied this to missing modality sentiment analysis, treating missing samples as augmented data. Li et al.(Li et al., 2022a) used MTD for Lidar-Radar segmentation, improving robustness against missing modalities. Self-distillation helps a model improve by learning from its own soft representations (Figure 12(b)). Hu et al. proposed the ShaSpec (Wang et al., 2023a), which utilizes distillation between modality-shared branches and modality-specific branches of available modalities to help improve model performance. Based on ShaSpec, researchers proposed Meta-learned Cross-modal Knowledge Distillation (Wang et al., 2024b) to further weigh the importance of different modalities to improve performance. Ramazanova et al. (Ramazanova et al., 2024) used mutual information and self-distillation for egocentric video tasks, making predictions invariant to missing modalities. Shi et al. proposed PASSION (Shi et al., 2024), a self-distillation method designed to use the multi-modal branch to help other uni-modal branches of avaiable modalities to improve multimodal medical segmentation performance with missing modality. Hybrid distillation methods combine various distillation approaches to improve student performance. For medical segmentation, Yang et al. (Yang et al., 2022) distilled teacher model knowledge at every decoder layer, outperforming ACNet(Wang et al., 2021). Wang et al. (Wang et al., 2023d) introduced ProtoKD, which captures inter-class feature relationships for improved segmentation under missing modality conditions. For sentiment analysis, CorrKD (Li et al., 2024a) leverages contrastive distillation and prototype learning to enhance performance in uncertain modality scenarios. Aforementioned methods effectively address the missing modality problem and achieve good generalization during testing. However, except for some intermediate and self-distillation methods, most assume the complete dataset is available for training (means teacher can access full-modality information during training), with missing modalities encountered only during testing. As a result, most are unsuitable for handling incomplete training datasets.
Figure 13. Illustration of graph fusion methods. We set modality 2 (white circle) as the missing modalities. Features from modality 1 and 3 can be aggregated via this graph fusion method and fused features maintains the same dimension.
2.3.3. Graph Learning Based Method
Graph-learning based methods leverage relationships between nodes and edges in graph-structured data for representation learning and prediction. We categorize approaches for addressing the missing modality problem into two main types: graph fusion and graph neural network (GNN) methods. Graph fusion methods integrate multimodal data using a graph structure (Figure 13), making them adaptable to various networks. For example, Angelou et al.(Angelou et al., 2019) proposed a method mapping each modality to a common space using graph techniques to preserve distances and internal structures. Chen et al.(Chen and Zhang, 2020) introduced HGMF, which builds complex relational networks using hyper edges that dynamically connect available modalities. Zhao et al.(Zhao et al., 2022) developed Modality-Adaptive Feature Interaction for Brain Tumor Segmentation, adjusting feature interactions across modalities based on binary existence codes. More recently, Yang et al.(Yang et al., 2023b) proposed a Graph Attention Based Fusion Block to adaptively fuse multimodal features, using attention-based message passing to share information between modalities. These fusion methods can be flexibly inserted into any network for integrating multiple modalities.
(a) Individual GNN Method
(b) Unified GNN Method
Figure 14. Illustration of graph learning based methods. We set modality 2 as the missing modalities. (a) Each modality first passes through DNN or builds its graph, and then they are uniformly processed using GNNs. (b) This method first complete the graph from missing-modality samples via approaches like FeatProp (Malitesta et al., 2024), then process it by Graph Neural Networks (GNN).
GNN methods directly encode multimodal information into a graph structure, with GNNs used to learn and fuse this information. Early approaches (Zhao et al., 2016; Wang et al., 2015) employed Laplacian graphs to connect complete and incomplete multimodal samples. Individual GNN methods (Figure 14(a)), such as DESAlign (Wang et al., 2024c), extract features using neural networks or GNNs and fuse them for prediction. Unified GNN methods ( Figure 14(b)) first complete the graph and then use GNNs for prediction, such as in Zhang et al.’s M3Care (Zhang et al., 2022), which uses adaptive weights to integrate information from similar patients. Lian et al.(Lian et al., 2023) proposed the Graph Completion Network, which reconstructs missing modalities by mapping features back to the input space. In recommendation systems, FeatProp(Malitesta et al., 2024) propagates known multimodal features to infer missing ones, while MUSE (Wu et al., 2024) represents patient-modality relationships as a bipartite graph and learns unified patient representations. In knowledge graphs, Chen et al. (Chen et al., 2023a) introduced Entity-Level Modality Alignment to dynamically assign lower weights to missing or uncertain modal information, reducing the risk of misguidance in learning. Above methods can leverage the graph structure to better capture relationships both within/between modalities and samples. However, these methods tend to have lower efficiency, scalability, and higher development complexity compared to other kind of approaches.
2.3.4. Multimodal Large Language Model
With the rise of LLMs like ChatGPT (Bubeck et al., 2023), their impressive generalization capabilities across tasks have demonstrated the transformative power. However, human understanding extends beyond language, incorporating modalities like vision and audio. This has led researchers to explore MLLMs, designed to handle diverse user inputs across modalities, including cases with missing modalities, leveraging the flexibility of Transformers. In current MLLM architectures, LLMs act as feature processors, integrating feature tokens from different modality-specific encoders and passing the output to task-/modality-specific decoders. This enables LLM to not only capture rich inter-modal dependencies, but also naturally carry the ability to handle any number of modalities, that is, the ability to solve the missing modality problem. Most MLLMs employ transformer-based modality encoders, such as CLIP, ImageBind (Girdhar et al., 2023), and LanguageBind (Zhu et al., 2023), which encode multimodal inputs into a unified representation space. Examples include BLIP-2 (Li et al., 2023b), which bridges the modality gap using a lightweight Querying Transformer, and LLaVA (Liu et al., 2024a), which enhances visual-language understanding with GPT-4-generated instruction data (Bubeck et al., 2023). These models are optimized for tasks like Visual Question Answering, Dialogue, and Captioning. Recent advancements extend output generation to multiple modalities, such as images. AnyGPT (Zhan et al., 2024) and NExT-GPT (Wu et al., 2023a) unify modalities like text, speech, and images using discrete representations and multimodal adaptors, enabling seamless multimodal interaction. CoDi (Tang et al., 2024) introduces Composable Multimodal Conditioning, allowing arbitrary modality generation through weighted feature summation. Other notable MLLMs include Qwen-VL (Bai et al., 2023), AnyMAL (Moon et al., 2023), and X-LLM (Chen et al., 2023b), showcasing the growing potential of MLLMs in diverse applications. Although MLLMs can flexibly handle any number of modalities, they have disadvantages of inconsistent multi-modal positional encoding, training difficulty, and high GPU resource requirements. Additionally, no specific MLLM benchmarks on missing modality problems have been found.
2.4. Model Selection
Model selection methods aim to use one or more selected models for downstream tasks, while also enhancing robustness and performance. These methods can be categorized into ensemble, dedicated, and discrete scheduler methods. Ensemble methods combine predictions from multiple selected models through, such as voting, weighted averaging, and similar approaches to improve the final decision-making accuracy and stability. Dedicated methods allocate different sub-tasks (e.g., different missing modality cases) to specialized individual models, focusing on specific sub-tasks or sub-datasets. In the discrete scheduler method, users can use natural language instructions to enable the LLMs to autonomously select the appropriate model based on types of modalities and downstream tasks.
(a) Model-Based Ensemble Method
(b) Modality-Based Ensemble Method
Figure 15. Two general architectures for ensemble methods. (a) The model-based ensemble methods are equipped with $n$ three-modal DNN models and obtain the final output based on the prediction results of all $n$ models. (b) In modality-based ensemble methods, a uni-modal DNN is equipped for each modality, and finally the results of these uni-modal DNNs are sent to ensemble methods.
2.4.1. Ensemble Method
Ensemble learning methods allow flexibility in supporting different numbers of expert models to combine their predictions, as shown in Figure 15. Early work (Wagner et al., 2011) in sentiment analysis employed ensemble learning to handle missing modalities, averaging predictions from modality-specific models. With deep learning advancements, the Ensemble-based Missing Modality Reconstruction network (Zeng et al., 2022) was introduced, leveraging weighted judgments from multiple full-modality models when generated missing modality features exhibit semantic inconsistencies. This type of model-based ensemble method, depicted in Figure 15(a), integrates various full-modality models to aid decision-making. Another type is modality-based ensemble methods, as shown in Figure 15(b), where each modality is paired with a uni-modal model, and only available modalities contribute to decision-making. In multimodal medical image diagnosis, early studies found uniformly weighted methods performed better than weighted averaging and voting approaches (Yuan et al., 2012). In multimodal object detection, Chen et al. (Chen et al., 2022) proposed a probabilistic ensemble method that flexibly handles missing modalities via probabilistic marginalization, proving highly efficient. Li et al. (Li et al., 2023a) recently proposed Uni-Modal Ensemble with Missing Modality Adaptation, training models per modality and performing post-fusion training. Other approaches (Miech et al., 2018) calculate feature-based weights to fuse modalities, where weights reflect feature importance for final predictions. A hybrid method, Dynamic Multimodal Fusion (DynMM) (Xue and Marculescu, 2023), uses gating functions to select uni-/multi-modal models dynamically. Lastly, Proximity-based Modality Ensemble (PME) (Cha et al., 2024) addresses negative fusion by employing a cross-attention mechanism with attention bias, adaptively reducing noise in multi-modal decoding.
Figure 16. General idea of Dedicated Methods. If modality 2 in the sample is missing. Dedicated Methods select the sample-suited model from DNN models library for training and testing. In above figure, the three modalities can form a total of seven models with different modality combinations. For easy understanding, we use colored circles in purple rectangles to represent the modalities that the model can handle. In order to adapt to the missing modality 2 (light green circles), Dedicated Methods usually select the DNN model (purple rectangle) that can handle modality 1 (light orange circles) and modality 3 (light blue circles) for training and testing.
2.4.2. Dedicated Method
Dedicated methods assign different tasks to specialized models. We show the general idea of the kind of method in Figure 16. KDNet (Hu et al., 2020) was the first dedicated method proposed to handle different combinations of missing modalities. It treats uni-modality specific models as student models, learning multimodal knowledge from the features and logits of a multimodal teacher model. Those trained uni-modals can be used for different missing modality cases. Following KDNet, ACNet (Wang et al., 2021) also utilizes this distillation method but introduces adversarial co-training, further improving the performance. Lee et al. (Lee et al., 2023b) proposed missing-aware prompts to address missing modality problems based on the prompt learning by using input- and attention- level prompts for each kind of missing modality case, only 1% of the model total parameters needed fine-tuning for downstream tasks. Similarly, some researchers (Reza et al., 2023) equip different adapter layers for each missing modality case to handle missing modality problems.
Figure 17. General idea of discrete scheduler methods. The user uses text instructions to make LLM call the downstream task modules to complete the processing of data samples and feedback to the user. As long as there are sufficiently diverse downstream task modules, it is possible to handle a variety of tasks with different numbers of modalities.
2.4.3. Discrete Scheduler Method
In discrete scheduler methods (Figure 17), LLMs act as task controllers, determining the execution order of different discrete sub-tasks broken down from the major task/instruction. While the LLM does not directly process multimodal data, it interprets language instructions and orchestrates task execution across uni- and multi-modal modules. This structured yet flexible approach is particularly effective for outputs requiring sequential tasks, enabling the system to handle any number of modalities and naturally addressing missing modality problems. For example, Visual ChatGPT (Wu et al., 2023b) integrates multiple foundation models with ChatGPT to enable interaction through text or/and images, allowing users to pose complex visual questions, provide editing instructions, and receive feedback within a multi-step collaboration framework. HuggingGPT (Shen et al., 2024) is an LLM-driven agent that manages and coordinates various AI models from Hugging Face. It leverages LLMs for task planning, model selection, and summarization to address complex multimodal tasks. ViperGPT (Surís et al., 2023) combines visual and language tasks into modular subprograms, generating and executing Python code for complex visual queries without additional training to achieve effective outputs. There are other similar discrete scheduler approaches, such as AudioGPT (Huang et al., 2024), MM-REACT (Yang et al., 2023a), and LLaVA-Plus (Liu et al., 2023a).
Some aforementioned dedicated and ensemble methods can flexibly handle the missing modality problem without additional training, but most of them require more model storage space, which is not feasible for many resource-constrained devices. For instance, DynMM requires storing various uni- and multi-modality models. As the number of modalities increases, the required number of models also rises. Also, modality-based ensemble methods struggle to adequately model the complex inter-modality relationships to make final predictions. In addition, discrete scheduler methods can solve a variety of tasks when there are sufficient types of downstream modules, but they usually require LLMs to respond quickly and understand human instructions accurately in real world scenarios.
2.5. Discussion
Table 1. Comparison of thirteen kinds of deep multimodal learning with missing modality methods from four dimensions.
Dimension Method Pros and Cons
Modality Augmentation Modality Composition Pros: Simple and effective for data augmentation. Cons: Unsuitable for pixel-level tasks (Composition); Modality number $\uparrow$, generative model number or training complexity $\uparrow$.
Modality Generation
Feature Space Engineering Regularization-based Pros: Flexible for novel modalities with better generalization. Cons: Hard to balance constraints and handle dataset imbalance.
Correlation Driven
Representation Composition
Representation Generation
Architecture Engineering Attention-based Pros: Attention methods scale well; Distillation and graph learning are effective. Cons: Most attention methods demand high computational resources; Distillation and graph learning struggle with incomplete datasets and training complexity.
Distillation-based
Graph Learning-based
Multimodal Large Language Model
Model Selection Ensemble Pros: Effective for specific tasks. Cons: Grows complex with more modalities; Suffers from modality imbalance; Demands significant storage space.
Dedicated
Discrete Scheduler
2.5.1. Summary
In this section, we categorize existing deep MLMM approaches across four dimensions: modality augmentation, feature space engineering, architecture engineering, and model selection. We further subdivide these into thirteen categories under a fine-grained methodology taxonomy. Table 1 summarizes the overall pros and cons of these methods from four dimensions. Generative and distillation methods are the most common approaches for addressing missing modalities; they are easy to implement and deliver strong performance. With the rise of Transformers, attention-based methods have become more popular due to their larger receptive fields and parallelism. However, indirect-to-task generation methods and most distillation methods (except for some intermediate and self-distillation methods) are currently unable to handle incomplete training datasets (means cannot access missing modality data during training). Below, we summarize the advantages and disadvantages of these thirteen methods from four dimensions.
Modality Augmentation: Modality composition methods operate directly on the original data level to address missing modality problems by combining existing data samples or padding missing data. These methods are straightforward and can serve as conventional data augmentation techniques. However, they are typically unsuitable for pixel-level downstream tasks and heavily rely on the existing modalities. On the other hand, modality generation methods employ generative models to synthesize missing modalities, mitigating the performance degradation caused by the missing data. Despite their advantages, these methods are often limited by the availability of full-modality samples in the dataset and the increased training complexity due to the additional modalities. They also require extra storage space for the generative models.
Feature Space Engineering: Regularization- and correlation-based representation constraints allow novel modalities to be introduced by simply adding corresponding branches. However, as the number of constraints increases, balancing them becomes challenging. Moreover, it is difficult to capture unique relationships across all modalities, especially when the dataset is highly imbalanced due to missing modalities. Correlation-based methods, in particular, often fail to handle missing modality samples in the test environment, as they are typically designed for pre-training with incomplete datasets. Similar to modality composition methods, representation composition methods attempt to recovery missing modality representations by combining representations of available modalities. However, since representations typically carry more generalized information than original data, these methods tend to yield better results. Representation generation methods further improve this process by generating representations based on relationships between modalities, allowing the reconstruction of missing modality representations. Many studies have confirmed the effectiveness of this approach in handling missing modalities. However, if the dataset inherently contains missing modalities, indirect-to-task representation generation methods that aim to recovery missing modality representations through the reconstruction of original data are not feasible. Furthermore, when the dataset has severely imbalanced modality combinations, generative methods may also fail, becoming overly reliant on existing modalities.
Architecture Engineering: Some methods focus on designing model training or inference architecture to alleviate the performance degradation caused by missing modalities. Attention-based methods are valued by researchers because they can effectively capture the relationships between various modalities, are scalable in terms of dataset size, and are highly parallelizable. Single attention mechanisms used for modality fusion have the advantage of being plug-and-play. However, the drawbacks of transformer-based methods in this category are also relatively obvious: training multimodal transformers from scratch requires excessive training time and many GPU resources. Recently popular PEL methods can effectively mitigate this drawback, but the generalization performance of this method is still not as well as full-parameter training or tuning. Although MLLMs can handle an arbitrary number of modalities with flexibility, they are constrained by training complexity and require substantial computational resources. Distillation-based methods are relatively easy to implement, with student models learning how to reconstruct missing modality representations and inter-modal relationships from teacher models. Since the teacher model typically receives full-modality samples as input, it can simulate any missing modality cases during training, ensuring strong performance. However, most distillation-based methods are limited to complete datasets and are inapplicable to datasets with inherent missing modalities. To our knowledge, only some intermediate distillation and self-distillation methods have attempted to input mismatched samples of the same class into the teacher for distillation purposes. Graph learning-based methods capture intra- and inter-modal relationships more effectively, but as the number of modalities increases, the development complexity and inefficiency of these methods become more pronounced. They are also currently unsuitable for large-scale datasets.
Model Selection: These methods select one or more models for prediction. Ensemble and dedicated methods are affected by imbalances in modality combinations within the dataset. As the number of modalities grows, the complexity and models to be trained also increase, especially for dedicated methods. The discrete scheduler method requires an LLM to coordinate, but if callable modules are insufficient, downstream tasks may not be completed. Additionally, inference speed is limited by the LLM and modules. Model selection methods also require significant model storage space.
2.5.2. Recovery and Non-Recovery Methods
Table 2. Another common taxonomy of recovery and non-recovery methods based on the three stages of early, intermediate and late in the conventional taxonomy of multimodal learning.
Methods Early Intermediate Late Hybrid
Recovery
Modality Composition (Yang et al., 2024)
Modality Generation (Wang et al., 2024a)
Representation Composition (Wang et al., 2023c)
Representation Generation (Woo et al., 2023)
Distillation-based (Shen and Gao, 2019)
Graph Learning-based (Malitesta et al., 2024)
Representation Generation (Ma et al., 2021a)
Distillation-based (Li et al., 2022a)
Correlation Driven (Ma et al., 2021b)
Distillation-based (Sikdar et al., 2023)
Non-Recovery Dedicated (Wang et al., 2021) Discrete Scheduler (Wu et al., 2023b) Representation Composition (Delgado-Escano et al., 2021) Attention-based (Ge et al., 2023) Graph Learning-based (Zhao et al., 2022) Multimodal Large Language Model (Wu et al., 2023a) Regularization-based (Liang et al., 2019) Distillation-based (Shi et al., 2024) Ensemble (Chen et al., 2022) /
Specifically, we have further classified our method categories in Table 1 into another common taxonomy to facilitate subsequent researchers to better distinguish, which is based on the three-stage (early, intermediate, and late) conventional taxonomy. In Table 2, we provide this taxonomy of our method categories dealing with missing modalities across these stages, based on whether they recover the missing modalities or not. We also attach a reference work for different methods. In our analysis of 315 papers, we found that 75.5% of the works focus on recovering missing modality information for inference, while only 24.6% explore inference without modality recovery. Among the recovery methods, early-stage and mid-stage modality recovery account for 20.3% and 45.8%, respectively, with late-stage and multi-stage recovery methods each accounting for 4.7%. For non-recovery methods, early, mid, and late-stage modality fusion represent 4.2%, 14.1%, and 6.3%, respectively. No non-recovery methods that combine multiple stages were found. We can observe that methods on recovering missing modality features in the middle stage account for the largest proportion. We think this is because recovering features, as opposed to the original data, can avoid much noise and bias while providing more modality-specific information. Additionally, compared to later-stage features, mid-stage features tend to be more enriched.
2.5.3. Limitation
Due to the variety of settings in most missing-modality works, it is difficult to compare the performance of different methods using specific dataset benchmarks. We leave this issue for future work.
3. Applications and Datasets
Over the past decade, multimodal learning based on deep learning has experienced explosive growth, leading to extensive applications in both academia and industry. Accompanying these applications is the emergence of various forms of multimodal datasets. However, the collection of such datasets is often labor-intensive and costly. In certain specific application directions, issues such as user privacy concerns, sensor malfunctions on data collection devices, and other factors can result in datasets with missing modalities. In severe cases, up to 90% of the samples may have missing modalities, making it challenging for conventional MLFM to achieve good performance during model training. This has given rise to the task of MLMM. Since the factors causing incomplete datasets usually stem from different application directions, we below introduce the corresponding datasets based on these applications: Sentiment Analysis, Medical Diagnosis, Retrieval/Captioning, Remote Sensing, Robotic Vision, and others. We also categorize those datasets according to the applications and data type in Table 3.
Table 3. Common and typical datasets used by deep MLMM methods divided by main applications and modality type. “Vision” includes data from visual sensors such as RGB, Depth, Infrared, LiDAR, Radar, Event, and Optical Flow. “Bio-Sensors” includes data from sensors like: electrophysiological and respiratory sensors.
Applications Modality Type Common and Typical Datasets
Sentiment Analysis Vision+Audio eNTERFACE’05 (Martin et al., 2006), RAVDESS (Livingstone and Russo, 2018), CREMA-D (Cao et al., 2014)
Vision+Text eBDtheque (Guérin et al., 2013), DCM (Nguyen et al., 2018), Manga 109 (Matsui et al., 2017)
Vision+Bio-Sensors Ulm-TSST (Stappen et al., 2021a, b), RECOLA (Ringeval et al., 2013), SEED (Zheng and Lu, 2015)
Vision+Audio+Text CMU-MOSI (Zadeh et al., 2016), CMU-MOSEI (Zadeh et al., 2018), IEMOCAP (Busso et al., 2008),MSP-IMPROV (Busso et al., 2016), MELD (Poria et al., 2018), CH-SIMI (Yu et al., 2020), ICT-MMMO (Wöllmer et al., 2013), YouTube (Morency et al., 2011)
Medical Diagnosis MRI, CT Scans BRATS (Menze et al., 2014; Myronenko, 2019), ADNI (Jack Jr et al., 2008), IXI [^1] Source: https://brain-development.org/ixi-dataset/
Bio-Sensors StressID (Chaptoukaev et al., 2024), PhysioNet (Kemp et al., 2000)
Electronic Health Record MIMIC-CXR (Johnson et al., 2019), MIMIC-IV (Johnson et al., 2023), NCH (Lee et al., 2022), BCNB (Xu et al., 2021), ODIR (Li et al., 2021)
Information Retrieval Vision+Text Amazon Series Datasets (Lakkaraju et al., 2013; McAuley et al., 2015; He and McAuley, 2016; PromptCloud, 2017), MIR-Flickr25K (Duin, [n. d.]), NUS-WIDE (Chua et al., 2009)
Vision+Audio+Text MSR-VTT (Xu et al., 2016), TikTok (Lin et al., 2023)
Remote Sensing Vision WHU-OPT-SAR (Li et al., 2022b), DFC2020 (Yokoya et al., 2020), MSAW (Shermeyer et al., 2020), Trento (Rasti et al., 2017), Houston (Piacentini, 2023), Berlin (Hong et al., 2021), GRSS (Debes et al., 2014)
Robotic Vision Vision
RegDB (Ye et al., 2021), SYSU-MM01 (Wu et al., 2017), UWA3DII (Rahmani et al., 2016),
Delivery (Zhang et al., 2023a), NuScenes (Caesar et al., 2020), MVSS (Ji et al., 2023)
Vision+Audio Ego4D-AR (Grauman et al., 2022), Epic-Sounds (Huh et al., 2023), Epic-Kitchens (Damen et al., 2022), Speaking Faces (John and Kawanishi, 2022)
Vision+Skeleton/Inertial Data MMG-Ego4D (Gong et al., 2023), Northwestern-UCLA (Wang et al., 2014), NTU60 (Shahroudy et al., 2016)
3.1. Sentiment Analysis
Typically, the goal of sentiment analysis is to classify the current emotional state by combining information from multiple modalities such as textual, auditory, and RGB information. This field has garnered significant attention from researchers due to its promising applications in various specific scenarios, including customer service, market research, health monitoring, entertainment, and advertising. In the early stages of audio-visual sentiment analysis research, researchers discovered that facial detection algorithm in data collection devices sometimes failed to capture faces (e.g., due to occlusion) or the recorded audio is too noisy to use (Cohen et al., 2004; Sebe et al., 2005), resulting in some data samples that only contained audio or image, leading to incomplete audio-visual datasets. Consequently, researchers began exploring methods to enable sentiment analysis models to continue functioning effectively when inputting data samples with missing modalities, and additionally incorporating textual data as an extra modality. Currently, the datasets for applying sentiment analysis can be roughly divided into two categories. One type of datasets  (Martin et al., 2006; Livingstone and Russo, 2018; Cao et al., 2014; Stappen et al., 2021a, b; Ringeval et al., 2013; Zheng and Lu, 2015; Zadeh et al., 2018; Busso et al., 2008, 2016; Poria et al., 2018; Yu et al., 2020; Wöllmer et al., 2013; Morency et al., 2011) involves using video, audio, text, and biosensors to determine human emotional states in real life or in movies. The other type (Guérin et al., 2013; Nguyen et al., 2018; Matsui et al., 2017) involves using text and images to assess emotional states in contexts such as comic books. Please refer to Table 3 for the classification based on modality type.
3.2. Medical Diagnosis
Medical diagnosis has always been the focus of attention of people all over the world, but the uneven distribution of medical resources, how to promote personalized medicine, improve medical efficiency and how to reduce diagnostic errors are all current pain points. The emergence of artificial intelligence is expected to improve these problems. Because medical diagnosis requires comprehensive judgment of information from various modalities such as medical history, physical examination and imaging data, which is exactly what multimodal learning is good at, most researchers are currently developing intelligent medical diagnosis systems based on multimodal learning. The datasets that current attempts used to solve the missing modality problem in multimodal medical diagnosis mainly focus on eight areas, such as Neuroimaging and Brain Disorders (Menze et al., 2014; Myronenko, 2019; Young et al., 2017; Jack Jr et al., 2008; Lesjak et al., 2018), Cardiovascular (Liu et al., 2016; Moody and Mark, 2001), Cancer Detection (Arya and Saha, 2020; Tomczak et al., 2015), Women Health Analysis (Zhang et al., 2023c; Xu et al., 2021), Ophthalmology (Li et al., 2021; Zhang et al., 2022), Sleep Disorders (Lee et al., 2022; Marcus et al., 2013), Clinical Predictions (Johnson et al., 2016, 2023, 2019), and Biomedical Analysis (Chaptoukaev et al., 2024; Schmidt et al., 2018; Kemp et al., 2000). We have categorized some commonly used and representative datasets from above by modality type, as shown in Table 3.
3.3. Information Retrieval
Information Retrieval is a technology that automatically retrieves relevant content or data based on queries, historical behavior, preferences, and attribute data. It has received widespread attention due to its success on various platforms such as search engines and social media. It improves user satisfaction and information access experience through algorithmic analysis and prediction of content that users may find relevant. Multimodal learning makes modern retrieval systems possible because it can effectively process and analyze multiple types of information including text, images, audio, etc. Due to user privacy concerns and data sparsity issues, some researchers have begun exploring MLMM approaches in retrieval systems. Many of these studies leverage datasets (Lakkaraju et al., 2013; McAuley et al., 2015; He and McAuley, 2016; PromptCloud, 2017; Duin, [n. d.]; Chua et al., 2009; Xu et al., 2016; Lin et al., 2023) from websites such as Amazon and TikTok for their research. See categories of those datasets based on modality type in Table 3.
3.4. Remote Sensing
By leveraging multimodal learning, we can integrate various types of visual data, such as Synthetic Aperture Radar data and multi-/hyper-spectral data. As a result, corresponding various datasets have also emerged (Yokoya et al., 2020; Shermeyer et al., 2020; Rasti et al., 2017; Piacentini, 2023; Hong et al., 2021; Debes et al., 2014). Those dataset enables the assessment and analysis of different environmental conditions, resources, and disasters on earth from satellites or aircraft. Such capabilities can significantly aid in environmental protection, resource management, and disaster response. In practical, in remote sensing tasks such as multimodal land cover classification tasks, optical modalities may be unavailable due to cloud cover or sensor damage. Therefore, it is also very necessary to solve the missing modality problem for remote sensing researchers.
3.5. Robotic Vision
Robotic vision is the technology that enables robots to perceive and understand their environment through visual sensors by acquiring and processing image data. This includes tasks such as object and facial recognition, environment modeling, navigation, and behavior understanding, allowing robots to autonomously perform complex operations and interact with humans. Typically, the modality combination formula in robotic vision tasks can be expressed as RGB+X, where X can include, but are not limited to, LiDAR, radar, infrared sensors, depth sensors and auditory sensors. Both RGB and other sensor data can be missing for various reasons. Below, we introduce five common tasks in robotic vision concerning the missing modality problems. (1) Multimodal Segmentation: This task aims to input multimodal data and use segmentation masks to locate objects of interest. In the context of missing modalities, this task is common in datasets for autonomous vehicles, where sensors might fail due to damage or adverse weather conditions. Typical data combinations include RGB+(Depth/Optical Flow/LiDAR/Radar/Infrared/Event data) (Takumi et al., 2017; Liu et al., 2023b; Caesar et al., 2020; Zhang et al., 2023a). Other datasets focus on indoor scene segmentation (Silberman et al., 2012; Ha et al., 2017) (RGB, Depth, Thermal) and material segmentation (Liang et al., 2022) (RGB, Angle of Linear Polarization, Degree of Linear Polarization, Near-Infrared images). (2) Multimodal Detection: Similar to multimodal segmentation, multimodal detection aims to locate objects of interest using bounding boxes. Commonly used multimodal detection combinations in the context of missing modalities include: RGB with Depth (Silberman et al., 2012; Lai et al., 2011) and RGB with Thermal (Hwang et al., 2015; kaggle, 2020). (3) Multimodal Activity Recognition: This task can input data from the aforementioned visual sensors and also use audio cues to recognize human activities. Common datasets (Grauman et al., 2022; Huh et al., 2023; Damen et al., 2022; Gong et al., 2023; Wang et al., 2014; Shahroudy et al., 2016) for this task are combinations of RGB with Audio, Depth, Thermal, Inertial and skeleton data. (4) Multimodal Person Re-Identification: This task aims to use depth and thermal (Ye et al., 2021; Wu et al., 2017) information to robustly identify individuals under varying lighting conditions, such as at night. (5) Multimodal Face Anti-Spoofing: This task utilizes multiple sensors (such as RGB, infrared, and depth cameras) and multiple types of information to detect and prevent spoofing in facial recognition systems (John and Kawanishi, 2022; Zhang et al., 2020; Liu et al., 2021b). Its purpose is to enhance the system’s ability to identify fake faces (such as photos, videos, masks, etc.) by comprehensively analyzing data from different modalities, thereby increasing the security and reliability of facial recognition systems.
3.6. Other Applications
Except above tasks, there are many other areas where MLMM methods are being explored. For example, conventional audio-visual classification (Vielzeuf et al., 2018); audio-visual question answering (Park et al., 2024); multimodal large language models in visual-dialogue (Bubeck et al., 2023), captioning (Liu et al., 2023a) and so on; hand pose estimation using depth images and heat distributions (Tompson et al., 2014; Choi et al., 2017); knowledge graph completion utilizing multimodal data with missing modalities (Toutanova et al., 2015; Liu et al., 2019); multimodal time series prediction, such as stock prediction (Luo et al., 2024) and air quality forecasting (Zheng et al., 2015); multimodal gesture generation on how to create natural animations (Liu et al., 2022); analyzing single-cell multimodal data (Mimitou et al., 2021; Swanson et al., 2021) in biology.
3.7. Discussion
Current MLMM research primarily focuses on the areas of sentiment analysis, medical diagnosis, remote sensing analysis, retrieval/captioning, and robotic vision. Among these, a significant portion of research is dedicated to addressing the missing modality problem in video, text, and audio-based sentiment analysis, MRI segmentation and clinical prediction in medical diagnosis, and multi-sensor semantic segmentation for autonomous vehicles. In contrast, research efforts in MLMM for streaming data and scientific fields are relatively limited. In addition, the majority of the publicly available datasets mentioned above are complete in terms of modalities, and naturally occurring datasets with missing modalities are rare. As a result, these tasks are often evaluated by considering all possible combinations of missing modalities based on the existing types of modalities in the dataset, and then performing training and testing on the performance of the missing modalities, supplemented by different missing modality rates. In the works we reviewed, 38% controlled for varying degrees of missing modality rates and validated accordingly, while the remaining 62% used random modality missing methods. When categorizing missing modality rates into mild (¡30%), moderate (30%-70%), and severe (¿70%) levels, the proportions are 72.6%, 82.2%, and 69.9%, respectively. It is worth noting that a single study may perform validation across multiple missing modality rate levels. Additionally, we found that 36.4% of the works trained on incomplete training datasets (cannot access missing modality data during training), while 63.6% actually focused on testing with missing modalities. We can see that there is still small number of research on incomplete training datasets and call on more researchers to pay attention to this field.
4. Open Issues and Future Research Directions
4.1. Debate Between Recovery and Non-Recovery Methods
In MLMM, “to recover or not to recover the missing modality”, that is also the question. According to our observations, most current approaches to handling MLMM can be broadly categorized into two types: one focuses on recovering the missing modal information so that the multimodal model can continue functioning (Wang et al., 2023a; Ma et al., 2021a), and the other attempts to make predictions using only the currently available modal information (Hu et al., 2020; Zhang et al., 2024). However, some studies (Yao et al., 2024) have shown that the recovered modal information might be ignored or dominated by the existing modal information due to the imbalance missing rates of different modalities in the training dataset, leading the model to still rely on the available modalities when making predictions. Moreover, other research (Lin et al., 2024) has found that in situations where crucial modal information is missing, the model might be unduly influenced by less important existing modal information, resulting in incorrect outcomes. In such cases, attempting to recover missing modal information can potentially alleviate this problem. Consequently, determining under what circumstances recovery or non-recovery methods are more beneficial for model predictions, as well as how to measure whether the recovered modal information is not dominated by other modal information or is unbiased, remains significant challenges. Addressing these issues is essential for advancing the effectiveness and reliability of deep multimodal learning with missing modality methods.
4.2. Efficiency and Lightweight
Current solutions for improving the performance of MLMM often overlook the need for more lightweight and efficient methods. For instance, some model selection methods (Hu et al., 2020; Xue and Marculescu, 2023) require training an independent model for each potential cases of missing modalities or a combination of many uni- and multi-modal models. Similarly, methods (Chen et al., 2024b; Zhang et al., 2023b) that aim to recover missing modal information typically involve using a distinct model for each kind of missing information or a large unified model to generate all modalities. Although these approaches generally perform well, they are often too heavy. In the real world, many multimodal models need to be deployed on resource-constrained devices, such as space or disaster-response robots, and mobile devices. These devices cannot accommodate high-performance GPUs and are prone to sensor damage, which can be difficult or costly to repair. Therefore, there is an urgent need to explore efficient and lightweight solutions for MLMM that can operate effectively under these constraints. Developing such solutions is critical for the practical deployment of MLMM models in real-world where resources are limited.
4.3. Multimodal Streaming Data
Currently, only part of research focuses on handling missing modalities in multimodal streaming data, mainly in sentiment analysis (Lin and Hu, 2023). However, long sequences of multimodal data, such as RGB+X video and multimodal time series data, are common in the real world and are essential for tasks like anomaly detection and video tracking on robots. Therefore, MLMM needs to address missing modality problem in streaming data to advance this field further.
4.4. Multimodal Reinforcement Learning
Multimodal reinforcement learning leverages information from different sensory modalities, enabling agents to learn effective strategies in environments. This approach has wide applications in areas such as robotic grasping, drone control, and driving decisions. However, these agent-based tasks often face practical challenges, as real-world scenarios frequently encounter sensor failures or restricted access to certain modal data, compromising the robustness of the algorithms. Currently, only a limited amount of work (Vasco et al., 2021; Mei et al., 2023; Lee et al., 2021; Chen et al., 2021; Awan et al., 2022) try to address the problem of missing data or modalities in reinforcement learning. Therefore, we hope the community can focus more on this practically significant missing modality problem in the reinforcement learning.
4.5. Multimodal AI for Natural Science
In scientific domains such as drug prediction (Deng et al., 2020), materials science (Muroga et al., 2023), and astronomy (Hong et al., 2023), the application of multimodal learning is of paramount importance. These fields inherently involve diverse types of data, from molecular structures and genomic sequences to spectral images and material properties. Multimodal learning holds the promise of integrating these varied data sources to enhance predictive accuracy and uncover novel insights. However, the practical implementation of multimodal learning in scientific research faces significant challenges. Data access restrictions, high acquisition costs, and the incompatibility of heterogeneous data sources often result in incomplete datasets or entirely missing modalities. These issues hinder the seamless integration of multimodal data, making it difficult to fully leverage the potential of multimodal learning in scientific experiments. Despite the critical need, there has been limited research (He et al., 2024) focusing on MLMM within the scientific field. Addressing this gap is crucial for advancing AI-driven scientific discoveries. Innovative approaches are required to handle incomplete multimodal dataset and develop robust models capable of performing well even when some modalities are absent. This will not only improve the applicability of AI in scientific research but also pave the way for breakthroughs in understanding and innovation.
4.6. Lack of Multimodal Large Language Model Benchmarks
The recent rise of large pre-trained models such as GPT4 has given researchers hope for achieving Artificial General Intelligence. Consequently, more and more large models (Touvron et al., 2023; Brown et al., 2020; Liu et al., 2024a) are integrating visual, auditory, and other modal information to realize the vision of Multimodal Large Language Models. Although some research has been done based on those models, a detailed analysis and benchmarks is urgently needed to evaluate the performance of large multimodal models on the missing modality problem.
5. Conclusion
In this survey, we present the first comprehensive review of Deep Multimodal Learning with Missing Modality. We begin with a brief introduction to the historical development of the missing modality problem and the real-world reasons that underscore its significance. Following this, we categorize and summarize the current advances of this field from two perspectives in a fine-grained way: methodologies, applications and datasets. Finally, we discuss the existing challenges and potential future directions in this field. Although more and more researchers are involved in studying the problem of missing modality, we are also concerned about some urgent issues that need to be addressed, such as a unified benchmark for testing (e.g., multimodal large language models) and the need for a wider range of applications (e.g., natural science). With our comprehensive and detailed approach, we hope that this survey will inspire more researchers to explore deep multimodal learning techniques with missing modalities, ultimately contributing to the development of robust and high-performance AI systems.